# Magnetic Polarisability of Octet Baryons via Lattice QCD

### v2

@@ II.2 Implementing the quark model
To present the predictions of the quark model for the magnetic polarisability as a function of quark mass, we require a relationship between the lattice baryon masses and the constituent quark masses. This relationship has been long established for the baryon octet at the experimental point and therefore we need only extend this relationship away from the physical point. In addition, the octet-decuplet mass splitting in Eq. II.1 demands knowledge of both octet and decuplet baryons. Similarly, $\expectationvalue{r^{2}}_{f}$ describing the distribution of quark flavours within the octet baryons is also required. We seek smooth, interpolated expressions for each of these quantities.
Our lattice calculations are performed on the PACS-CS gauge-field ensembles Aoki et al. (2009) and therefore we draw on their published baryon masses in the Sommer scheme. Ref. Walker-Loud (2014) demonstrated that the nucleon masses observed in finite-volume lattice QCD simulations display a ruler-style linear behaviour when plotted as a function of $m_{\pi}$. This provides a simple characterisation of lattice baryon masses and in the spirit of the simple quark model, we take this approach to interpolate the PACS-CS baryon masses for the octet and decuplet.
To estimate the constituent quark masses away from the physical point, we utilise a simple linear model
@@
which leverages the maximum difference in strange quarks in the baryon octet. Here $m^{0}_{\rm oct}$ and $\alpha_{\rm oct}$ are fit parameters, and $m_{l}$ and $m_{h}$ are the light and heavy constituent quark masses already determined in the simple constituent quark model summarised in Ref. Workman et al. (2022). The parameter $m^{0}_{\rm oct}$ allows some of the nucleon mass to have its origin in the confining potential and $\alpha_{\rm oct}$ allows for spin-dependent effects to change the slope from 1. Using experimental baryon masses and the constituent quark masses $m_{l}=338\,$MeV and $m_{h}=510\,$MeV from the PDG Workman et al. (2022), we solve for the fit parameters $m^{0}_{\rm oct}$ and $\alpha_{\rm oct}$.
@@
The fit parameters corresponding to the two models are shown in Tab. 1. We see that in the octet case the slope parameter $\alpha_{\rm oct}=1.09$ is larger than unity. This is in accord with anticipated hyperfine effects. Because the strength of the attractive hyperfine interaction is inversely proportional to the product of the constituent quark masses, the attraction diminishes with increasing quark mass. Thus the baryon mass grows at a rate exceeding the rate of the quark mass increase.
@@
Table 1: Fit parameters for the simple models of Eq. II.2 and Eq. II.2.
@@
We use the octet fit parameters and the ruler-style interpolated $N$ and $\Xi$ masses to obtain the constituent quark masses as a function of $m_{\pi}$. These masses are given in Tab. 2. We see the expected approach to the physical constituent quark mass 338 MeV Workman and Others (2022) while the strange quark mass remains stable as expected.
Table 2: Constituent quark masses for the PACS-CS gauge field ensembles Aoki et al. (2009). Constituent masses are obtained from the model Eq. II.2 with $(\alpha_{\rm oct},m^{0}_{\rm oct})=$($1.09,-169\,$MeV). All masses are in MeV.
kappa $m_{\pi}^{\rm PACS-CS}$ $m_{l}$ $m_{h}$
@@
Referring back to Eq. II.1 for the magnetic polarisability, the intrinsic quark magnetic moments contained within the octet-decuplet baryon magnetic transition matrix element are simply
@@
Quark distribution radii on the PACS-CS ensembles have been determined by Stokes et al. Stokes et al. (2020) where they examined the proton and neutron. We interpolate the proton and neutron squared radii as linear in $\log(m_{\pi})$ and this provides an excellent description of those lattice results. Eq. II.1 requires quark distribution radii for single quark flavours of unit charge, as the charge factors appear elsewhere in the expression. We define radii for the doubly, $D$, and singly, $S$, represented quark sectors for single quark flavours of unit charge by
@@
$\expectationvalue{r^{2}}_{D}$ $=\frac{1}{2}\left(2\expectationvalue{r^{2}}_{p}-\expectationvalue{r^{2}}_{n}\right),$ ((33))
$\expectationvalue{r^{2}}_{S}$ $=\expectationvalue{r^{2}}_{p}-2\expectationvalue{r^{2}}_{n}.$ ((34))
We approximate the doubly and singly represented strange-quark distribution radii to be that of the light quark at the heaviest PACS-CS quark mass where the light quark mass is similar to that of the strange quark as seen in  Tab. 2.
Figure 1: Quark model predictions for the magnetic polarisabilities of octet baryons are plotted as a function of the squared pion mass as a proxy for the quark mass. Dashed lines represent the pion masses of the PACS-CS ensembles. The legend is ordered to match the vertical ordering at the physical point.
With the octet and decuplet baryon masses, constituent quark masses, magnetic moments, and quark flavour distribution radii parameterised as a function of $m_{\pi}^{2}$, we are now able to present values for the magnetic polarisability as predicted by the constituent quark model as illustrated in Fig. 1. We note that the predicted magnitudes of the proton and neutron magnetic polarisabilities are somewhat larger than expected from experimental measurements. While we will address this issue in the following, the trends of the model are worthy of discussion.
First, we note the quark model predicts qualitatively different values for the negatively charged baryons. Their polarisabilities are predicted to be opposite in sign and much smaller in magnitude. Recalling Eq. II.1 with the transition magnetic moment written explicitly in terms of quark moments using equation Eq. (23)
@@
we first highlight the squared-difference between the magnetic moments of each sector
@@
When the quark charges of the sectors are opposed, such as in the proton, these quark moments sum to produce a large contribution. Conversely, in the case of the negative baryons $\Xi^{-}\,$and $\Sigma^{-}\,$, the $d$ and $s$ quarks both carry a charge of $-1/3$ providing cancellation between the terms. However, the splitting of the $d$ and $s$ quark masses admits a small contribution. It is this cancellation that gives rise to the very small nature of the magnetic polarisabilities of the negative baryons. Further cancellation comes from the opposite sign contribution of $\beta_{1}$ and $\beta_{2}$. It’s clear the $\beta_{2}$ term describing quark distributions dominates the sum.
Conversely, it is the opposition of quark sector electric charges that is the main generator of large magnitudes for the magnetic polarisability. Moreover, this is achieved through the inclusion of a $u$ quark whose charge magnitude is double that of the $d$ or $s$ quarks. When added and squared the contribution of opposing charge pairs, $u-d$ or $u-s$, is an order of magnitude larger than the contribution of a single $d$ or $s$ quark flavour.
@@
Given that the quark distribution radius, $\expectationvalue{r^{2}}_{f}$, is also smaller for the strange quark, together these effects suppresses the magnitude of the contributions of the strange quark.
@@
As a result, the polarisability is very sensitive to the up quark. The up quark generates a large $\beta_{1}$, but having two up quarks causes a large opposite contribution from $\beta_{2}$. This explains why the $\Xi^{0}\,$and $n$ have greater magnetic polarisabilities than the $\Sigma^{+}\,$and proton respectively.
@@
Another observation is that the mass splitting between the octet and decuplet baryons in $\beta_{1}$ benefits the hyperon magnetic polarisabilities over the nucleon polarisabilities. In the nucleon, the scalar diquark is composed of two light quarks. The strength of this hyperfine interaction is inversely proportional to the product of the constituent quark masses; hence the hyperfine attraction is strong in the nucleon and the hyperfine repulsion is strong in the $\Delta$. Together, these hyperfine interactions produce a large mass splitting in the magnetic transition term, $\beta_{1}$. In the hyperons considered herein, the scalar diquark is composed of a light $u$ or $d$ quark and a strange quark. This time the magnitude of the hyperfine interaction is weaker and the octet-decuplet mass splitting is smaller. Thus the magnitude of the magnetic transition term, $\beta_{1}$, is enhanced. This effect explains why the magnitude of the $\Xi^{0}\,$polarisability exceeds that of the neutron and would suggest the $\Sigma^{+}\,$exceeds the proton.
@@
The alternate ordering of the magnetic polarisability of the proton and $\Sigma^{+}\,$highlights the complexity of the magnetic polarisability. Here the larger mass of the strange quark causes a reduction in the transition magnetic moment which outweighs that of the octet-decuplet mass splitting and the reduced magnitude of $\beta_{2}$.
@@
The $\Xi^{0}\,$and $\Sigma^{+}\,$both have a $u$-$s$ scalar diquark and therefore $\beta_{1}$ is similar. The difference in the total polarisability is associated with $\beta_{2}$ where the second $u$ quark in $\Sigma^{+}\,$makes a larger negative contribution to the polarisability of the $\Sigma^{+}\,$compared to the second $s$ quark in the $\Xi^{0}\,$whose magnitude is small due to both the charge and the mass of the strange quark.
@@
Finally, we note that the magnetic polarisability of the $\Xi^{0}\,$rises quite steeply as the pion mass approaches the physical point. For all baryons, both $\beta_{1}$ and $\beta_{2}$ increase in magnitude as the physical pion mass is approached. However, for the $\Xi^{0}\,$the increase in $\beta_{1}$ significantly outpaces the increase in $\beta_{2}$ near the physical point. It is a combination of all aspects discussed thus far that make this possible. The magnetic transition term, $\beta_{1}$ benefits from both the presence of a $u$ quark enhancing the numerator, but also the presence of a $u$-$s$ scalar diquark suppressing the mass splitting and thus further enhancing $\beta_{1}$. Moreover, the presence of two $s$ quarks ensures $\beta_{2}$ is a small as possible, through the small electric charge, the suppression factor of the strange quark mass and the reduced distribution radius.
@@ III Background Field Implementation
The background field method is a well established approach to calculating the magnetic polarisability in a lattice QCD calculation. The discussion here is based on Ref. Primer et al. (2014) founded on Ref. Smit and Vink (1987).
@@
This relation is exact for a constant background field as all higher order terms involve a second or higher order derivative. The plaquette at coordinates $x,y$
@@
giving the desired field away from the $y$-direction boundary. However, on a finite lattice with sites labelled $(0\leq x/a\leq N_{x}-1)$ and $(0\leq y/a\leq N_{y}-1)$, there is a discontinuity at the $y$-boundary due to the periodic boundary conditions used for the QCD fields. To address this, we make use of the first term from Eq. (40), $\partial_{x}A_{y}$ and assign
@@
This ensures that the discontinuity at the boundary in the $y$-direction at $y/a=N_{y}-1$ is compensated via the $y$-boundary contribution from $A_{y}$.
Having used both terms available in Eq. (40), the double boundary, $x/a=N_{x}-1$ and $y/a=N_{y}-1$, gives rise to a quantisation condition for the field strength. Here the plaquette takes the value
@@
which provides the required value when the second exponential $\exp(-i\,a^{2}\,N_{x}\,N_{y}\,q_{f}\,e\,B)=1$. Setting $a^{2}\,N_{x}\,N_{y}\,q_{f}\,e\,B$ to be an integer multiple of $2\pi$ provides the field-strength quantisation
@@
Throughout this work, we will specify the field quanta in terms of the charge of the down quark. The quantisation condition becomes
@@
While in principle the fermion propagator source can be placed anywhere within the spatial volume, one obtains the optimal signal to noise in the baryon correlation functions when the source is placed at the origin of the electromagnetic potential Primer et al. (2014). Hence as one increases statistics by considering many spatial source positions on a single gauge field, we cycle the selected fermion source to the (0,0,0) position by periodic circular shifts of the gauge field. Then the magnetic field is introduced as a phase on the gauge field links as described above.
@@ IV Lattice QCD Formalism
In the presence of a uniform background magnetic field, the energy of a baryon changes as a function of magnetic field strength Martinelli et al. (1982); Primer et al. (2014)
@@
Here the mass of the baryon, $m$, is complemented by contributions from the magnetic moment $\vec{\mu}$, the Landau term proportional to $|q_{B}\,e\,B|$ where $q_{B}$ is the charge of the baryon, and the magnetic polarisability $\beta$. For neutral baryons, the Landau term does not contribute, and for charged baryons, we will use a U(1) Landau-mode projection to select $n=0$. This aspect of the calculation is detailed in subsection V.4.
This expression arises as a non-relativistic Taylor expansion of the relativistic energy Itzykson and Zuber (2006). The large mass scale of the octet baryons enables the use of the non-relativistic approximation with systematic errors of ($\order{1\%}$ ) at the field strengths considered in this work. See for example, Chapter 3.6.2 of Ref. Bignell (2020). This approach simplifies the required calculations and improves the signal relevant to the polarisabilities.
The magnetic polarisability may be extracted by isolating the quadratic $B^{2}$ term in Eq. (47). The sign of the magnetic moment term is dependent on the alignment of the baryon’s spin, quantised in the $z$-direction, and the magnetic field. Hence it may be eliminated by summing energies associated with different spin orientations
@@
The magnetic polarisability and Landau terms are now isolated as the only remaining field strength dependent terms. This process could be mirrored in a lattice QCD calculation, determining the effective energies in the aligned and anti-aligned case, however this neglects the opportunity to cancel the highly correlated QCD fluctuations contained within the correlation functions of different field strengths and spin alignments.
To optimise this cancellation, we define a spin-field aligned correlator
@@
where the baryon’s spin is aligned with the magnetic field and a spin-field anti-aligned correlator
@@
As such, we define the magnetic polarisability energy shift $\delta E_{\beta}(B,t)$
@@
We note that this is the analogue to Eq. (48). The first term, the Landau term, vanishes for neutral baryons allowing for direct access to the magnetic polarisability. For charged baryons, the Landau term must be carefully considered in the fitting process.
Fitting a single-parameter quadratic fit to Eq. IV with the Landau term fully specified as discussed in Sec. VI allows the extraction of the magnetic polarisability.

### v3

@@ II.2 Implementing the quark model
To present the predictions of the quark model for the magnetic polarizability as a function of quark mass, we require a relationship between the lattice baryon masses and the constituent quark masses. This relationship has been long established for the baryon octet at the experimental point and we need only extend this relationship away from the physical point. In addition, the octet-decuplet mass splitting in Eq. II.1 demands knowledge of both octet and decuplet baryons. Similarly, $\expectationvalue{r^{2}}_{f}$ describing the distribution of quark flavors within the octet baryons is also required. We seek smooth, interpolated expressions for each of these quantities.
Our lattice calculations are performed on the PACS-CS gauge-field ensembles Aoki et al. (2009) and therefore we draw on their published baryon masses in the Sommer scheme Sommer (1994). Reference Walker-Loud (2014) demonstrated that the nucleon masses observed in finite-volume lattice QCD simulations display a ruler-style linear behavior when plotted as a function of $m_{\pi}$. This provides a simple characterization of lattice baryon masses and in the spirit of the simple quark model, we take this approach to interpolate the PACS-CS baryon masses for the octet and decuplet.
To estimate the constituent quark masses away from the physical point, we utilize a simple linear model
@@
which leverages the maximum difference in strange quarks in the baryon octet. Here $m^{0}_{\rm oct}$ and $\alpha_{\rm oct}$ are fit parameters, and $m_{l}$ and $m_{h}$ are the light and heavy constituent quark masses already determined in the simple constituent quark model summarized in Ref. Workman et al. (2022). The parameter $m^{0}_{\rm oct}$ allows some of the nucleon mass to have its origin in the confining potential and $\alpha_{\rm oct}$ allows for spin-dependent effects to change the slope from 1. Using experimental baryon masses and the constituent quark masses $m_{l}=338$ and $m_{h}=510\,$MeV from the PDG Workman et al. (2022), we solve for the fit parameters $m^{0}_{\rm oct}$ and $\alpha_{\rm oct}$.
@@
The fit parameters corresponding to the two models are shown in Table 1. We see that in the octet case the slope parameter $\alpha_{\rm oct}=1.09$ is larger than unity. This is in accord with anticipated hyperfine effects. Because the strength of the attractive hyperfine interaction is inversely proportional to the product of the constituent quark masses, the attraction diminishes with increasing quark mass. Thus the baryon mass grows at a rate exceeding the rate of the quark mass increase.
@@
Table 1: Fit parameters for the simple models of Eqs. II.2 and II.2.
@@
We use the octet fit parameters and the ruler-style interpolated $N$ and $\Xi$ masses to obtain the constituent quark masses as a function of $m_{\pi}$. These masses are given in Table 2. We see the expected approach to the physical constituent quark mass of 338 MeV Workman and Others (2022) while the strange-quark mass remains stable as expected.
Table 2: Constituent quark masses for the PACS-CS gauge-field ensembles Aoki et al. (2009). Constituent masses are obtained from the model Eq. II.2 with $(\alpha_{\rm oct},m^{0}_{\rm oct})=$($1.09,-169\,$MeV). All masses are in MeV.
$\kappa$ $m_{\pi}^{\rm PACS-CS}$ $m_{l}$ $m_{h}$
@@
Referring back to Eq. II.1 for the magnetic polarizability, the intrinsic quark magnetic moments contained within the octet-decuplet baryon magnetic transition matrix element are simply
@@
Quark distribution radii on the PACS-CS ensembles have been determined by Stokes et al. Stokes et al. (2020) where they examined the proton and neutron. We interpolate the proton and neutron squared radii as linear in $\log(m_{\pi})$ and this provides an excellent description of those lattice results. Equation. II.1 requires quark distribution radii for single quark flavors of unit charge, as the charge factors appear elsewhere in the expression. We define radii for the doubly $D$ and singly $S$ represented quark sectors for single quark flavors of unit charge by
@@
$\expectationvalue{r^{2}}_{D}$ $=\frac{1}{2}\left(2\expectationvalue{r^{2}}_{p}+\expectationvalue{r^{2}}_{n}\right),$ ((33))
$\expectationvalue{r^{2}}_{S}$ $=\expectationvalue{r^{2}}_{p}+2\expectationvalue{r^{2}}_{n}.$ ((34))
We approximate the doubly and singly represented strange-quark distribution radii to be that of the light-quark at the heaviest PACS-CS quark mass where the light quark mass is similar to that of the strange quark as seen in  Table 2.
Figure 1: Quark model predictions for the magnetic polarizabilities of octet baryons are plotted as a function of the squared pion mass as a proxy for the quark mass. Dashed lines represent the pion masses of the PACS-CS ensembles. The legend is ordered to match the vertical ordering at the physical point.
With the octet and decuplet baryon masses, constituent quark masses, magnetic moments, and quark flavor distribution radii parameterized as a function of $m_{\pi}^{2}$, we are now able to present values for the magnetic polarizability as predicted by the constituent quark model as illustrated in Fig. 1. We note that the predicted magnitudes of the proton and neutron magnetic polarizabilities are somewhat larger than expected from experimental measurements. While we will address this issue in the following, the trends of the model are worthy of discussion.
First, we note the quark model predicts qualitatively different values for the negatively charged baryons. Their polarizabilities are predicted to be opposite in sign and much smaller in magnitude. Recalling Eq. II.1 with the transition magnetic moment written explicitly in terms of quark moments using equation Eq. (23)
@@
we first highlight the squared difference between the magnetic moments of each sector
@@
When the quark charges of the sectors are opposed, such as in the proton, these quark moments sum to produce a large contribution. Conversely, in the case of the negative baryons $\Xi^{-}\,$and $\Sigma^{-}\,$, the $d$ and $s$ quarks both carry a charge of $-1/3$ providing cancellation between the terms. However, the splitting of the $d$ and $s$ quark masses admits a small contribution. It is this cancellation that gives rise to the very small nature of the magnetic polarizabilities of the negative baryons. Further cancellation comes from the opposite sign contribution of $\beta_{1}$ and $\beta_{2}$. It is clear the $\beta_{2}$ term describing quark distributions dominates the sum.
Conversely, it is the opposition of quark sector electric charges that is the main generator of large magnitudes for the magnetic polarizability. Moreover, this is achieved through the inclusion of a $u$ quark whose charge magnitude is double that of the $d$ or $s$ quarks. When added and squared the contribution of opposing charge pairs, $u-d$ or $u-s$, is an order of magnitude larger than the contribution of a single $d$ or $s$ quark flavor.
@@
Given that the quark distribution radius $\expectationvalue{r^{2}}_{f}$ is also smaller for the strange quark, together these effects suppress the magnitude of the contributions of the strange quark.
@@
As a result, the polarizability is very sensitive to the up quark. The up quark generates a large $\beta_{1}$, but having two up quarks causes a large opposite contribution from $\beta_{2}$. This explains why the $\Xi^{0}\,$and $n$ have greater magnetic polarizabilities than the $\Sigma^{+}\,$and proton respectively.
@@
Another observation is that the mass splitting between the octet and decuplet baryons in $\beta_{1}$ benefits the hyperon magnetic polarizabilities over the nucleon polarizabilities. In the nucleon, the scalar diquark is composed of two light quarks. The strength of this hyperfine interaction is inversely proportional to the product of the constituent quark masses; hence the hyperfine attraction is strong in the nucleon and the hyperfine repulsion is strong in the $\Delta$. Together, these hyperfine interactions produce a large mass splitting in the magnetic transition term, $\beta_{1}$. In the hyperons considered herein, the scalar diquark is composed of a light $u$ or $d$ quark and a strange quark. This time the magnitude of the hyperfine interaction is weaker and the octet-decuplet mass splitting is smaller. Thus the magnitude of the magnetic transition term, $\beta_{1}$, is enhanced. This effect explains why the magnitude of the $\Xi^{0}\,$polarizability exceeds that of the neutron and would suggest the $\Sigma^{+}\,$exceeds the proton.
@@
The alternate ordering of the magnetic polarizability of the proton and $\Sigma^{+}\,$highlights the complexity of the magnetic polarizability. Here the larger mass of the strange quark causes a reduction in the transition magnetic moment which outweighs that of the octet-decuplet mass splitting and the reduced magnitude of $\beta_{2}$.
@@
The $\Xi^{0}\,$and $\Sigma^{+}\,$both have a $u$-$s$ scalar diquark and therefore $\beta_{1}$ is similar. The difference in the total polarizability is associated with $\beta_{2}$ where the second $u$ quark in $\Sigma^{+}\,$makes a larger negative contribution to the polarizability of the $\Sigma^{+}\,$compared to the second $s$ quark in the $\Xi^{0}\,$whose magnitude is small due to both the charge and the mass of the strange quark.
@@
Finally, we note that the magnetic polarizability of the $\Xi^{0}\,$rises quite steeply as the pion mass approaches the physical point. For all baryons, both $\beta_{1}$ and $\beta_{2}$ increase in magnitude as the physical pion mass is approached. However, for the $\Xi^{0}\,$the increase in $\beta_{1}$ significantly outpaces the increase in $\beta_{2}$ near the physical point. It is a combination of all aspects discussed thus far that make this possible. The magnetic transition term, $\beta_{1}$ benefits from both the presence of a $u$ quark enhancing the numerator, but also the presence of a $u$-$s$ scalar diquark suppressing the mass splitting and thus further enhancing $\beta_{1}$. Moreover, the presence of two $s$ quarks ensures $\beta_{2}$ is as small as possible, through the small electric charge, the suppression factor of the strange-quark mass and the reduced distribution radius.
@@ III Background Field Implementation
The background-field method is a well-established approach to calculating the magnetic polarizability in a lattice QCD calculation. The discussion here is based on Ref. Primer et al. (2014) founded on Ref. Smit and Vink (1987).
@@
This relation is exact for a constant background field as all higher-order terms involve a second- or higher-order derivative. The plaquette at coordinates $x,y$,
@@
giving the desired field away from the $y$-direction boundary. However, on a finite lattice with sites labeled $(0\leq x/a\leq N_{x}-1)$ and $(0\leq y/a\leq N_{y}-1)$, there is a discontinuity at the $y$ boundary due to the periodic boundary conditions used for the QCD fields. To address this, we make use of the first term from Eq. (40), $\partial_{x}A_{y}$ and assign
@@
This ensures that the discontinuity at the boundary in the $y$-direction at $y/a=N_{y}-1$ is compensated via the $y$ boundary contribution from $A_{y}$.
Having used both terms available in Eq. (40), the double boundary, $x/a=N_{x}-1$ and $y/a=N_{y}-1$, gives rise to a quantization condition for the field strength. Here the plaquette takes the value
@@
which provides the required value when the second exponential $\exp(-i\,a^{2}\,N_{x}\,N_{y}\,q_{f}\,e\,B)=1$. Setting $a^{2}\,N_{x}\,N_{y}\,q_{f}\,e\,B$ to be an integer multiple of $2\pi$ provides the field-strength quantization
@@
Throughout this work, we will specify the field quanta in terms of the charge of the down quark. The quantization condition becomes
@@
While in principle the fermion propagator source can be placed anywhere within the spatial volume, one obtains the optimal signal-to-noise ratio in the baryon correlation functions when the source is placed at the origin of the electromagnetic potential Primer et al. (2014). Hence, as one increases statistics by considering many spatial source positions on a single gauge field, we cycle the selected fermion source to the (0,0,0) position by periodic circular shifts of the gauge field. Then the magnetic field is introduced as a phase on the gauge-field links as described above.
@@ IV Lattice QCD Formalism
In the presence of a uniform background magnetic field, the energy of a baryon changes as a function of magnetic-field strength Martinelli et al. (1982); Primer et al. (2014)
@@
Here the mass of the baryon, $m$, is complemented by contributions from the magnetic moment $\vec{\mu}$, the Landau term proportional to $|q_{B}\,e\,B|$ where $q_{B}$ is the charge of the baryon, and the magnetic polarizability $\beta$. For neutral baryons, the Landau term does not contribute, and for charged baryons, we will use a U(1) Landau-mode projection to select $n=0$. This aspect of the calculation is detailed in subsection V.4.
This expression arises as a nonrelativistic Taylor expansion of the relativistic energy Itzykson and Zuber (2006). The large mass scale of the octet baryons enables the use of the nonrelativistic approximation with systematic errors of ($\order{1\%}$ ) at the field strengths considered in this work. See for example, Chap. 3.6.2 of Ref. Bignell (2020). This approach simplifies the required calculations and improves the signal relevant to the polarizabilities.
The magnetic polarizability may be extracted by isolating the quadratic $B^{2}$ term in Eq. (47). The sign of the magnetic moment term is dependent on the alignment of the baryon’s spin, quantized in the $z$-direction, and the magnetic field. Hence, it may be eliminated by summing energies associated with different spin orientations
@@
The magnetic polarizability and Landau terms are now isolated as the only remaining field-strength-dependent terms. This process could be mirrored in a lattice QCD calculation, determining the effective energies in the aligned and antialigned case, however this neglects the opportunity to cancel the highly correlated QCD fluctuations contained within the correlation functions of different field strengths and spin alignments.
To optimize this cancellation, we define a ”spin-field aligned” correlator
@@
where the baryon’s spin is aligned with the magnetic field and a ”spin-field antialigned” correlator
@@
As such, we define the magnetic polarizability energy shift $\delta E_{\beta}(B,t)$
@@
We note that this is the analog to Eq. (48). The first term, the Landau term, vanishes for neutral baryons allowing for direct access to the magnetic polarizability. For charged baryons, the Landau term must be carefully considered in the fitting process.
Fitting a single-parameter quadratic fit to Eq. IV with the Landau term fully specified as discussed in Sec. VI allows the extraction of the magnetic polarizability.